When a rotating disk is observed from a standstill at some point, it spins 15 rounds at 5s and then 24 turns. Find the number of rotations in another 5s and the number of rotations of the disk from stillness at the moment of initial observation.


1
Expert's answer
2021-06-18T09:45:46-0400

Gives

Turns (n)=24

15round is time=5sec

1round is time="\frac{5}{15}"

24round is time ="\frac{5}{15}\times24=8sec"

"w=2\pi n=\frac{2\pi}{8}\rightarrow(1)"

Turn=24

T=10sec

Number of rotation

"w=\frac{2\pi x}{10}\rightarrow(2)"

Where x=rotation

Equation (1)and(2)are equal

"\frac{2\pi}{8}=\frac{2\pi x}{10}"

"x=\frac{10}{8}=1.25"

Number turn=24

Total time t=10sec rotation

"x=1.25\times24=30"